OpenStudy (anonymous):

r=1 -1 2 2 find the point which transform to point : 3 2 under T

OpenStudy (anonymous):

any body to help

OpenStudy (anonymous):

you will want to row reduce the augmented matrix: 1 -1 3 2 2 2 This will give you something in the form of: 1 0 a 0 1 b the point (a,b) will be your answer?

OpenStudy (anonymous):

er...i dont know how that question mark got there <.< its not supposed to be there lol. that will give you your answer.

OpenStudy (anonymous):

The question is this : Given that T is a linear transformation such that T= 1 -1 2 2 find the piont which is transformed onto the point 3 2 under T

OpenStudy (anonymous):

right, i know thats the question. what i posted is how to solve it.

myininaya (myininaya):

2(1 -1 3 ) 1 -1 3 1 -1 3 2 2 2 = 4 0 2 = 1 0 1/2 =1 0 1/2 = -1( 1 0 1/2) = 1 0 1/2 1 -1 3 1 -1 3 0 -1 3 =1 0 1/2 0 1 -3

myininaya (myininaya):

made a mistake

myininaya (myininaya):

2(1 -1 3 ) 1 -1 3 1 -1 3 2 2 2 = 4 0 8 = 1 0 2 =1 0 2 = -1( 1 0 2) = 1 0 2 1 -1 3 1 -1 3 0 -1 1 =1 0 2 0 1 1

OpenStudy (anonymous):

what is the answer